Circular No. 6720

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

MICROLENSING EVENT A. C. Becker, D. P. Bennett, N. Butler, B. A. Peterson, S. H. Rhie, S.

E. Thomson, S. E. Turner, and the MACHO collaboration (cf. IAUC 6312) write: “We predict a high-magnification for event 97-BLG-56, a gravitational microlensing event in progress toward the Galactic Bulge. The source star, located at α = 17^{h}55^{m}28^{s}.7, δ = −29^{o}33′42′′ (equinox 2000.0), appears to be a giant with unamplified magnitudes of V = 16.1, R = 14.8. The current microlensing fit predicts a peak amplification of > 5 on 1997 Aug.

24. Due to the star type and the high magnification of this event, it is expected that finite source effects such as changes in color and spectral-line equivalent widths will be observable. Spectroscopy and infrared photometry at or near peak will be invaluable in studying this unusual event. A finding chart and additional information are available by anonymous

ftp at darkstar.astro.washington.edu:/macho/Alert/97-BLG-56 or on
the WWW at http://darkstar.astro.washington.edu/. Potential ob-

servers are asked to email gman@mso.anu.edu.au or telephone E. Turner at (Australia) 6-249-0209 or 6-288-4727 to coordinate observations.”

AG DRACONIS M. Mikolajewski, Centre for Astronomy, N. Copernicus University,

Torun, reports an increase in brightness since April in all optical bands: Apr. 26.90 UT, U = 11.26, B = 11.31, V = 9.91, R = 8.44, I = 7.78; Aug. 16.95, 9.27, 9.99, 9.08, 7.78, 7.27; 17.85, 9.21, 9.95, 9.09, 7.77, 7.27; 18.86, 9.17, 9.94, 9.05, 7.73, 7.21.

E. Waagen, AAVSO, reports that, after being at minimum for ∼ 7.5 yr, AG Dra brightened in May–July 1994 (to m_{v} = 8.3), then again in June– July 1995 (mag 8.9) and June–July 1996 (mag 9.4). The current brightening began in late May 1997. Recent visual magnitude estimates: Aug. 1.93 UT, 9.1 (G. Poyner, Birmingham, England); 5.46, 9.2 (G. Hanson, Cave Creek, AZ); 15.94, 8.8 (Poyner); 17.96, 9.2 (B. Granslo, Fjellhamar, Norway).

V4368 SAGITTARII G. J. Garradd, Loomberah, N.S.W., reports the following precise CCD

position for this object (cf. IAUC 5961, 5976), obtained on Aug. 19.395 UT (mag R = 10.5): α = 18^{h}54^{m}40^{s}.34, δ = −19^{o}41′59′′.7 (equinox 2000.0).

Recent AAVSO visual magnitude estimates: June 28.2 UT, 10.5 (J. Bortle, Stormville, NY); July 13.8, 10.2 (L. A. G. Monard, Pretoria, South Africa); Aug. 2.8, 10.2 (Monard); 6.8, 10.3 (Monard); 10.8, 10.2 (Monard).

1997 August 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green